OpenStudy (anonymous):

When the coordinates (1, 1), (2, 3), (5, 3), and (4, 1) are joined, which shape is formed? Parallelogram Rectangle Rhombus Square

OpenStudy (anonymous):

@arabpride

OpenStudy (anonymous):

parallelogram

OpenStudy (anonymous):

have you done a really fast sketch of the dots? it ends up being a parallelogram.

OpenStudy (anonymous):

Parallelogram (Note NOT all sides are equal in length)

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

shorter sides are = sqrt(2^2 + 1^2) = sqrt(5) in length and longer sides are 3 in length